# A JWST Survey of the Supernova Remnant Cassiopeia A

### v1

@@ A JWST Survey of the Supernova Remnant Cassiopeia A
We present initial results from a JWST survey of the youngest Galactic core-collapse supernova remnant Cassiopeia A (Cas A), made up of NIRCam and MIRI imaging mosaics that map emission from the main shell, interior, and surrounding circumstellar/interstellar material (CSM/ISM). We also present four exploratory positions of MIRI/MRS IFU spectroscopy that sample ejecta, CSM, and associated dust from representative shocked and unshocked regions. Surprising discoveries include: 1) a web-like network of unshocked ejecta filaments resolved to 0.01 pc scales exhibiting an overall morphology consistent with turbulent mixing of cool, low-entropy matter from the progenitor’s oxygen layer with hot, neutrino and radioactively heated high-entropy matter, 2) a thick sheet of dust-dominated emission from shocked CSM seen in projection toward the remnant’s interior pockmarked with small () round holes formed by 01 knots of high-velocity ejecta that have pierced through the CSM and driven expanding tangential shocks, 3) dozens of light echoes with angular sizes between to reflecting previously unseen fine-scale structure in the ISM. NIRCam observations place new upper limits on infrared emission ( nJy at 3 m) from the neutron star in Cas A’s center and tightly constrain scenarios involving a possible fallback disk. These JWST survey data and initial findings help address unresolved questions about massive star explosions that have broad implications for the formation and evolution of stellar populations, the metal and dust enrichment of galaxies, and the origin of compact remnant objects.
@@ 1 Introduction
| Instrument | Filter | BW | PSF | Sources of strong emission | ||
| (m) | (m) | (″) | (sec) | |||
| NIRCam | F162M | 1.626 | 0.168 | 0.055 | 3350 | [Fe ii] 1.644; [Si i] 1.645; synchrotron |
| F356W | 3.563 | 0.787 | 0.116 | 1675 | [Ca iv] 3.207, [Si ix] 3.936; PAHs; synchrotron; dust | |
| F444W | 4.421 | 1.024 | 0.145 | 1675 | [Si ix] 3.936; [Ca v] 4.159; [Mg iv] 4.487, [Ar vi] 4.530; | |
| [K iii] 4.618; CO; synchrotron; dust | ||||||
| MIRI | F560W | 5.6 | 1.2 | 0.207 | 1598 | [Mg v] 5.61; dust; synchrotron |
| F770W | 7.7 | 2.2 | 0.269 | 1598 | [Ar ii] 6.99; PAHs, dust | |
| F1000W | 10.0 | 2.0 | 0.328 | 1598 | [Ar iii] 8.991; [S iv] 10.511; dust | |
| F1130W | 11.3 | 0.7 | 0.375 | 1598 | PAHs; dust | |
| F1280W | 12.8 | 2.4 | 0.420 | 1598 | [Ne ii] 12.814; [Ne V] 14.32; dust | |
| F1800W | 18.0 | 3.0 | 0.591 | 1598 | [Fe ii] 17.94; [S iii] 18.713; dust; H | |
| F2100W | 21.0 | 5.0 | 0.674 | 1598 | [S iii] 18.713; dust | |
| F2550W | 25.5 | 4.0 | 0.803 | 1598 | [O iv] 25.89; dust |
@@
Cassiopeia A (Cas A) is arguably the SNR that provides the clearest access to the properties of a core-collapse SN (Koo & Park, 2017). Cas A is the youngest Galactic core-collapse SNR known ( yr; Fesen et al. 2006c); it is among the closest ( kpc; Reed et al. 1995; Alarie et al. 2014); it is the only core-collapse remnant with a secure SN classification from light echo spectroscopy performed from multiple lines of sight (Type IIb; Krause et al. 2008; Rest et al. 2008, 2011); it is one of the best case studies to understand dust formation in SN ejecta and shock-processing of that dust (Rho et al., 2009; De Looze et al., 2017); and its central X-ray point source is a key object to understanding NS evolution models (Pavlov & Luna, 2009; Gotthelf et al., 2013; Posselt & Pavlov, 2022; Shternin et al., 2023). Models for the remnant suggest that the 15–25 M zero-age-main-sequence progenitor star lost the majority of its mass prior to explosion as a –6 M star (Chevalier & Oishi, 2003; Hwang & Laming, 2012; Lee et al., 2014), which was likely encouraged through interaction with a binary companion (Young et al., 2006; Sato et al., 2020). Although claims have been made of surviving OB companions in extragalactic SNe IIb (Maund et al., 2004; Ryder et al., 2018), HST observations have ruled out this possibility for Cas A and to date no surviving companion has been located (Kochanek, 2018; Kerzendorf et al., 2019).
@@ 2 Observations
Cas A was observed with JWST in Cycle 1 General Observers (GO) Program 1947 (PI: Milisavljevic). The observations reported here were carried out between August and November 2022, using NIRCam (Rieke et al., 2023) and MIRI (Wright et al., 2023). NIRSpec (Jakobsen et al., 2022) observations obtained as part of this program that overlap with the MIRI/MRS positions are reported in De Looze et al. (2024) and Rho et al. (2024). All JWST data used in this paper can be found in MAST: 10.17909/szf2-bg42.
@@
Unfortunately, there was an error with the tiling pattern of the spacecraft, which introduced gaps in the mosaics and a gradual east-to-west drift between the centers of the fields of view of approximately 16. We applied and were approved for a return visit to cover the gaps with three additional MIRI pointings per filter. These return observations were performed on 2022 October 26. The overlapping field center of all mosaics is approximately (J2000) = 23232697, (J2000) = +58086, with the entire FOV spanning approximately and rotated with position angle 131.5. This return visit provided a serendipitous opportunity to study variable light echoes in the region surrounding Cas A (see Section 5).
The imaging data were processed using the calibration pipeline version 1.8.4 and the calibration reference data system (CRDS) version 11.16.14, with the CRDS context file . Background images were produced from dedicated sky observations and subtracted from the individual images in the pipeline step. We did not subtract background images from the F770W, F1000W, F1130W, and F1280W data due to the variable spatial structure of the sky background at these wavelengths. The level 2 images were astrometrically aligned using the JWST Alignment Tool (JHAT) (Rest et al., 2023).
Mosaic images were then constructed using the default pipeline parameters but with the and steps turned off. Pixel scales of the final mosaics are 0.031 per pixel for the NIRCam SW camera images, and 0.063 for the NIRCam LW camera images. The pixel scale of the MIRI mosaics is 0.111. Composite images made from the mosaics are shown in Figure 1. Mosaics of individual filters showing the entire fields of view are shown in the Appendix.
@@
Below we present and briefly discuss all four MRS spectra, shown in Figure 2, with locations identified in Figure 3. We focus on P2 and P4 within the context of investigating unshocked interior ejecta. De Looze et al. (2024) provide an in-depth analysis of P2 within the context of interaction between the remnant and the dusty circumstellar environment (see also section 3), and Rho et al. (2024) investigate P1 and P3 within the context of molecule formation and destruction.
@@ 3.1 NIRCam
The strongest emission in the NIRCam mosaic (Figure 1) is seen along the main shell of ejecta that has encountered the reverse shock, and represents the remnant’s densest material ( cm) with temperatures between 5000 to 10,000 K (Chevalier & Kirshner, 1978; Hurford & Fesen, 1996; Smith et al., 2009; Lee et al., 2017). As seen in HST data (Fesen et al., 2001), the ejecta are often grouped into large filament complexes of varying scales, and are often increasingly dissipated in directions away from the reverse shock (Morse et al., 2004).
The superb resolution of the NIRCam/F162M (0.055) compared to earlier HST NIR images places a new constraint on the clump size as small as A.U. Furthermore, NIRCam observations detect numerous ejecta knots beyond the main ejecta shell, many of which have not been previously observed either in HST optical or ground-based NIR observations (Fesen & Milisavljevic, 2016; Koo et al., 2018). Some knots are visible only in the F162M image, while others are visible in both the F162M and F444W images, suggesting their elemental compositions are different. The physical and chemical properties of the outlying ejecta knots will be investigated by Koo et al. (2024).
@@
Clumpy N- and He-rich CSM (aka “quasi-stationary flocculi” or QSFs; see van den Bergh 1971, Koo et al. 2018, Koo et al. 2020) appear bright in the F162M image due to their strong [Fe ii] lines. This is evident because all prominent QSFs identified in the 2013 ground-based [Fe ii] 1.64 m image by Koo et al. (2018) remain observable in the new JWST data. JWST’s resolution exposes intricate morphologies arising from interactions with the SNR shock. Some QSFs have sharp boundaries with abrupt brightness drops implying the presence of unshocked CSM that has not been processed by the shock yet. Scattered throughout the interior are numerous never-before-seen complete and partial rings approximately 1 in size, associated with emission from shocked CSM, that overlap with holes observed in MIRI images (these features are discussed more extensively below in section 3.2).
@@ 3.2 MIRI
One possible interpretation of the Green Monster and its unusual pockmarked morphology is that it is associated with unshocked ejecta that have been expanded by small clumps of radioactive material (Ni or Ti). Cas A is already known to have been strongly shaped by large-scale Ni-bubbles that expanded into O-, Ar-, and S-rich ejecta, forming large-scale rings as observed in the shocked main shell (Lawrence et al., 1995; Blondin et al., 2001; DeLaney et al., 2010; Milisavljevic & Fesen, 2013; Alarie et al., 2014), and cavities as observed in the interior unshocked ejecta (Milisavljevic & Fesen, 2015), with radii of 0.5 to 1. Hence, the holes in the Green Monster could be a much smaller-scale version of this phenomenon, induced by extremely finely clumped radioactive material, expanding within the central interior of the remnant.
However, in light of low velocity ( km s) emission lines of [Fe ii] 1.644 m, H Br 4.05 m, [Ne ii] 12.81 m, and [Ne iii] 15.56 m associated with the small rings (De Looze et al., 2024), along with the X-ray properties of emission in this region that include correlation with morphology observed in the infrared (Vink et al. 2024; see also Section 6), we instead favor that the dominant emission in the Green Monster region is due to circumstellar gas, rich in dust, that has been excited by the forward shock and sculpted by stellar debris. In this scenario, the central location within the main shell is simply a projection effect, and the holes are the result of small knots ( 01) of high-velocity ejecta that have punctured through the CSM and driven expanding tangential shocks (e.g., Orlando et al. 2022).
Finally, an apparent blowout protruding outward from the main shell is seen in the western portion of the remnant (Figure 3, panel 5). This blowout runs due west and is different from the direction of the known southwest outflow (or “jet”) (Fesen, 2001; Hwang et al., 2004). The morphology of this western blowout shares some resemblance to the NE rupture in the main shell (Figure 3, panel 3; see also Fesen & Gunderson 1996), and the general location is known to host a large population of QSFs (Koo et al., 2018). The blowout could be associated with interaction between the remnant and an abrupt density discontinuity in the ISM, not unlike the breakout observed in the NE section of the Galactic remnant CTB 1 (Fesen et al., 1997). Nonetheless, the MIRI emission in this region has conspicuous overlap with i) Si-enriched X-ray emitting ejecta (Vink, 2004) and ii) the projected trajectories of dozens of high-velocity ( km s) S-rich ejecta knots moving radially outward from the center of expansion (Fesen & Milisavljevic, 2016), which together indicate that the blowout, like the NE rupture, may reflect structure imprinted by the explosion dynamics of the original SN (Laming et al., 2006).
@@ 4 IFU Spectroscopy
Our MIRI/MRS spectra only cover four very small fields of view (channel dependent, ranging from 3.5″ to 7″111See the MIRI MRS page in the JWST Documentation for details, https://jwst-docs.stsci.edu/jwst-mid-infrared-instrument/miri-observing-modes/miri-medium-resolution-spectroscopy), but have far superior spatial and spectra resolution compared to previous low-resolution m spectral maps of Cas A obtained by Spitzer with the Infrared Spectrograph (IRS) (Smith et al., 2009). An overview of the spectra of all four positions is shown in Figure 2. These MIRI/MRS positions sample a diverse range of emission from the remnant and surrounding environment. P1 and P3 were selected as regions covering different elemental compositions of ejecta and dust excited by the reverse shock in the north and south; both the shape of the dust continuum and the relative intensity of the emission lines vary between the two positions. P2 is coincident with one of the small holes of the Green Monster, but also samples a column of unshocked ejecta and dust through the center of the remnant. P4 was a “deep drilling” IFU position designed to be primarily sensitive to unshocked ejecta, especially iron, through the center of the remnant. Selection of these areas was guided in part by the Spitzer IRS spectral maps. Based on the in-flight performance report from Argyriou et al. (2023), MIRI/MRS wavelength resolution is accurate to 2-27 km s depending on wavelength, and has a spectrophotometric precision of 5.6%.

### v2

@@ A JWST Survey of the Supernova Remnant Cassiopeia A
We present initial results from a JWST survey of the youngest Galactic core-collapse supernova remnant Cassiopeia A (Cas A), made up of NIRCam and MIRI imaging mosaics that map emission from the main shell, interior, and surrounding circumstellar/interstellar material (CSM/ISM). We also present four exploratory positions of MIRI/MRS IFU spectroscopy that sample ejecta, CSM, and associated dust from representative shocked and unshocked regions. Surprising discoveries include: 1) a web-like network of unshocked ejecta filaments resolved to 0.01 pc scales exhibiting an overall morphology consistent with turbulent mixing of cool, low-entropy matter from the progenitor’s oxygen layer with hot, high-entropy matter heated by neutrino interactions and radioactivity, 2) a thick sheet of dust-dominated emission from shocked CSM seen in projection toward the remnant’s interior pockmarked with small () round holes formed by 01 knots of high-velocity ejecta that have pierced through the CSM and driven expanding tangential shocks, 3) dozens of light echoes with angular sizes between to reflecting previously unseen fine-scale structure in the ISM. NIRCam observations place new upper limits on infrared emission ( nJy at 3 m) from the neutron star in Cas A’s center and tightly constrain scenarios involving a possible fallback disk. These JWST survey data and initial findings help address unresolved questions about massive star explosions that have broad implications for the formation and evolution of stellar populations, the metal and dust enrichment of galaxies, and the origin of compact remnant objects.
@@ 1 Introduction
| Instrument | Filter | BW | PSF | Date | Sources of strong emission | ||
| (m) | (m) | (″) | (sec) | (UT) | |||
| NIRCam | F162M | 1.626 | 0.168 | 0.055 | 3350 | 2022 Nov 5 | [Fe II] 1.644; [Si I] 1.645; synchrotron |
| F356W | 3.563 | 0.787 | 0.116 | 1675 | 2022 Nov 5 | [Ca IV] 3.207; [Si IX] 3.936; PAHs; synchrotron; dust | |
| F444W | 4.421 | 1.024 | 0.145 | 1675 | 2022 Nov 5 | [Si IX] 3.936; [Ca V] 4.159; [Mg IV] 4.487; | |
| [Ar VI] 4.530; [K III] 4.618; CO; synchrotron; dust | |||||||
| MIRI | F560W | 5.6 | 1.2 | 0.207 | 1598 | 2022 Aug 4-5, Oct 22 | [Mg V] 5.61; dust; synchrotron |
| F770W | 7.7 | 2.2 | 0.269 | 1598 | 2022 Aug 4-5, Oct 22 | [Ar II] 6.99; PAHs, dust | |
| F1000W | 10.0 | 2.0 | 0.328 | 1598 | 2022 Aug 4-5, Oct 22 | [Ar III] 8.991; [S IV] 10.511; dust | |
| F1130W | 11.3 | 0.7 | 0.375 | 1598 | 2022 Aug 4-5, Oct 22 | PAHs; dust | |
| F1280W | 12.8 | 2.4 | 0.420 | 1598 | 2022 Aug 4-5, Oct 22 | [Ne II] 12.814; [Ne V] 14.32; dust | |
| F1800W | 18.0 | 3.0 | 0.591 | 1598 | 2022 Aug 4-5, Oct 22 | [Fe II] 17.94; [S III] 18.713; dust; H2 | |
| F2100W | 21.0 | 5.0 | 0.674 | 1598 | 2022 Aug 4-5, Oct 22 | [S III] 18.713; dust | |
| F2550W | 25.5 | 4.0 | 0.803 | 1598 | 2022 Aug 4-5, Oct 22 | [O IV] 25.89; dust |
@@
Cassiopeia A (Cas A) is arguably the SNR that provides the clearest access to the properties of a core-collapse SN (Koo & Park, 2017). Cas A is the youngest Galactic core-collapse SNR known ( yr; Fesen et al. 2006c); it is among the closest ( kpc; Reed et al. 1995; Alarie et al. 2014); it is the only core-collapse remnant with a secure SN classification from light echo spectroscopy performed from multiple lines of sight (Type IIb; Krause et al. 2008; Rest et al. 2008, 2011); it is one of the best case studies to understand dust formation in SN ejecta and shock-processing of that dust (Rho et al., 2009; De Looze et al., 2017); and its central X-ray point source is a key object to understanding NS evolution models (Pavlov & Luna, 2009; Gotthelf et al., 2013; Posselt & Pavlov, 2022; Shternin et al., 2023). Models for the remnant suggest that the 15–25 M⊙ zero-age-main-sequence progenitor star lost the majority of its mass prior to explosion as a –6 M⊙ star (Chevalier & Oishi, 2003; Hwang & Laming, 2012; Lee et al., 2014), which was likely encouraged through interaction with a binary companion (Young et al., 2006; Sato et al., 2020). Although claims have been made of surviving OB companions in extragalactic SNe IIb (Maund et al., 2004; Ryder et al., 2018), HST observations have ruled out this possibility for Cas A and to date no surviving companion has been located (Kochanek, 2018; Kerzendorf et al., 2019).
@@ 2 Observations
Cas A was observed with JWST in Cycle 1 General Observers (GO) Program 1947 (PI: Milisavljevic). The observations reported here were carried out between August and November 2022, using NIRCam (Rieke et al., 2023) and MIRI (Wright et al., 2023). NIRSpec (Jakobsen et al., 2022) observations obtained as part of this program that overlap with the MIRI/MRS positions are reported in De Looze et al. (in prep.) and Rho et al. (2024). All JWST data used in this paper can be found in MAST: 10.17909/szf2-bg42.
@@
Unfortunately, there was an error with the tiling pattern of the spacecraft, which introduced gaps in the mosaics and a gradual east-to-west drift between the centers of the fields of view of approximately 16′′. We applied and were approved for a return visit to cover the gaps with three additional MIRI pointings per filter. These return observations were conducted on 2022 October 26 and performed with a position angle approximately 90 degrees to the original visit, which made the edges of the detectors parallel. The overlapping field center of all mosaics is approximately (J2000) = 23h23m2697, (J2000) = +58o086, with the entire FOV spanning approximately and rotated with position angle 131.5.
The imaging data were processed using the JWST Science Calibration Pipeline (Bushouse et al., 2022) version 1.8.4 and the calibration reference data system (CRDS) version 11.16.14, with the CRDS context file . Background images were produced from dedicated sky observations and subtracted from the individual images in the pipeline step. We did not subtract background images from the F770W, F1000W, F1130W, and F1280W data due to the variable spatial structure of the sky background at these wavelengths. The level 2 images were astrometrically aligned using the JWST Alignment Tool (JHAT) (Rest et al., 2023).
Mosaic images were then constructed using the default pipeline parameters but with the and steps turned off. Pixel scales of the final mosaics are 0.031′′ per pixel for the NIRCam SW camera images, and 0.063′′ for the NIRCam LW camera images. The pixel scale of the MIRI mosaics is 0.111′′. Composite images made from the mosaics are shown in Figure 1. Mosaics of individual filters showing the entire fields of view are shown in the Appendix.
@@
Below we present and briefly discuss all four MRS spectra, shown in Figure 2, with locations identified in Figure 3. We focus on P2 and P4 within the context of investigating unshocked interior ejecta. De Looze et al. (in prep.) provide an in-depth analysis of P2 within the context of interaction between the remnant and the dusty circumstellar environment (see also section 3), and Rho et al. (2024) investigate P1 and P3 within the context of molecule formation and destruction.
@@ 3.1 NIRCam
The strongest emission in the NIRCam mosaic (Figure 1) is seen along the main shell of ejecta that has encountered the reverse shock, and represents the remnant’s densest material ( cm-3) with temperatures between 5000 to 10,000 K (Chevalier & Kirshner, 1978; Hurford & Fesen, 1996; Smith et al., 2009; Lee et al., 2017). As seen in HST data (Fesen et al., 2001), the ejecta are often grouped into large filament complexes of varying scales, and are often increasingly dissipated in directions away from the reverse shock (Morse et al., 2004).
The superb resolution of the NIRCam/F162M (0.055′′) compared to earlier HST NIR images places a new constraint on the clump size as small as A.U. Furthermore, NIRCam observations detect numerous ejecta knots beyond the main ejecta shell, many of which have not been previously observed either in HST optical or ground-based NIR observations (Fesen & Milisavljevic, 2016; Koo et al., 2018). The knots show different colors in the NIRCam three-color image in Figure 1, suggesting their elemental compositions are different. The physical and chemical properties of the outlying ejecta knots will be investigated by Koo et al. (in prep.).
@@
Clumpy N- and He-rich CSM (aka “quasi-stationary flocculi” or QSFs; see van den Bergh 1971, Koo et al. 2018, Koo et al. 2020) appear bright in the F162M image due to their strong [Fe II] lines. This is evident because all prominent QSFs identified in the 2013 ground-based [Fe II] 1.64 m image by Koo et al. (2018) remain observable in the new JWST data. JWST’s resolution exposes intricate morphologies arising from interactions with the SNR shock. Some QSFs have sharp boundaries with abrupt brightness drops implying the presence of unshocked CSM that has not been processed by the shock yet. Scattered throughout the interior are numerous never-before-seen complete and partial rings approximately 1′′ in size, associated with emission from shocked CSM, that overlap with holes observed in MIRI images (these features are discussed more extensively below in section 3.2).
@@ 3.2 MIRI
One possible interpretation of the Green Monster and its unusual pockmarked morphology is that it is associated with unshocked ejecta that have been expanded by small clumps of radioactive material (56Ni or 44Ti). Cas A is already known to have been strongly shaped by large-scale Ni-bubbles that expanded into O-, Ar-, and S-rich ejecta, forming large-scale rings as observed in the shocked main shell (Lawrence et al., 1995; Blondin et al., 2001; DeLaney et al., 2010; Milisavljevic & Fesen, 2013; Alarie et al., 2014), and cavities as observed in the interior unshocked ejecta (Milisavljevic & Fesen, 2015), with radii of 0.5′ to 1′. Hence, the holes in the Green Monster could be a much smaller-scale version of this phenomenon, induced by extremely finely clumped radioactive material, expanding within the central interior of the remnant.
However, in light of low velocity ( km s-1) emission lines of [Fe II] 1.644 m, H Br 4.05 m, [Ne II] 12.81 m, and [Ne III] 15.56 m associated with the small rings (De Looze et al. in preparation), along with the X-ray properties of emission in this region that include correlation with morphology observed in the infrared (Vink et al. 2024; see also Section 6), we instead favor that the dominant emission in the Green Monster region is due to circumstellar gas, rich in dust, that has been excited by the forward shock and sculpted by stellar debris. In this scenario, the central location within the main shell is simply a projection effect, and the holes are the result of small knots ( 01) of high-velocity ejecta that have punctured through the CSM and driven expanding tangential shocks (e.g., Orlando et al. 2022).
Finally, an apparent blowout protruding outward from the main shell is seen in the western portion of the remnant (Figure 3, panel 5). This blowout runs due west and is different from the direction of the known southwest outflow (or “jet”) (Fesen, 2001; Hwang et al., 2004). The morphology of this western blowout shares some resemblance to the NE rupture in the main shell (Figure 3, panel 3; see also Fesen & Gunderson 1996), and the general location is known to host a large population of QSFs (Koo et al., 2018). The blowout could be associated with interaction between the remnant and an abrupt density discontinuity in the ISM, not unlike the breakout observed in the NE section of the Galactic remnant CTB 1 (Fesen et al., 1997). Nonetheless, the MIRI emission in this region has conspicuous overlap with i) Si-enriched X-ray emitting ejecta (Vink, 2004) and ii) the projected trajectories of dozens of high-velocity ( km s-1) S-rich ejecta knots moving radially outward from the center of expansion (Fesen & Milisavljevic, 2016), which together indicate that the blowout, like the NE rupture, may reflect structure imprinted by the explosion dynamics of the original SN (Laming et al., 2006).
@@ 4 IFU Spectroscopy
Our MIRI/MRS spectra only cover four very small fields of view (channel dependent, ranging from 3.5″ to 7″111See the MIRI MRS page in the JWST Documentation for details, https://jwst-docs.stsci.edu/jwst-mid-infrared-instrument/miri-observing-modes/miri-medium-resolution-spectroscopy), but have far superior spatial and spectra resolution compared to previous low-resolution m spectral maps of Cas A obtained by Spitzer with the Infrared Spectrograph (IRS) (Smith et al., 2009). An overview of the spectra of all four positions is shown in Figure 2. These MIRI/MRS positions sample a diverse range of emission from the remnant and surrounding environment. P1 and P3 were selected as regions covering different elemental compositions of ejecta and dust excited by the reverse shock in the north and south; both the shape of the dust continuum and the relative intensity of the emission lines vary between the two positions. P2 is coincident with one of the small holes of the Green Monster, but also samples a column of unshocked ejecta and dust through the center of the remnant. P4 was a “deep drilling” IFU position designed to be primarily sensitive to unshocked ejecta, especially iron, through the center of the remnant. Selection of these areas was guided in part by the Spitzer IRS spectral maps. Based on the in-flight performance report from Argyriou et al. (2023), MIRI/MRS wavelength resolution is accurate to 2-27 km s-1 depending on wavelength, and has a spectrophotometric precision of 5.6%.